OpenStudy (anonymous):

if s^m=64 and s and m are both integers,what is the least possible value of s-m? a)4 b)-4 c)-6 d)-8 e)-12 ?????

OpenStudy (anonymous):

hmm i can make it -10, so i guess it must be -12, but i don't see how. if i put \[s=-8, m=2\] then \[(-8)^2=64\] and \[s-m=-8-2=-10\]

OpenStudy (anonymous):

i am gonna die man